OpenStudy (anonymous):

if f(x)= x-2 / x+1 , then f(n+1) equals what?

OpenStudy (amistre64):

f(n+1) = (n+1) - 2 ------- (n+1) + 1

OpenStudy (amistre64):

might wanna simplify it...

OpenStudy (anonymous):

\[(n(n+3)/(n+1)\]

OpenStudy (amistre64):

??? you sure bout that ???

OpenStudy (anonymous):

i chose n+2-2/ n+1+1 and it was wrong so i think it is the other one

OpenStudy (amistre64):

n+2-2 ?? that aint right (n+1) - 2

OpenStudy (anonymous):

Just simply replace x with n f(n) = n-2/n+1 f(n+1)= n-1/n+2

OpenStudy (anonymous):

Sorry. I used the incorrect function syntax in a program solver. \[((n+1)-2)/((n+1)+1)=(n-1)/(n+2)\] amistre64 and sstarcia are correct.